# Benign overfitting in leaky ReLU networks with moderate input dimension

### v1

@@ Abstract
The problem of benign overfitting asks whether it is possible for a model to perfectly fit noisy training data and still generalize well. We study benign overfitting in two-layer leaky ReLU networks trained with the hinge loss on a binary classification task. We consider input data which can be decomposed into the sum of a common signal and a random noise component, which lie on subspaces orthogonal to one another. We characterize conditions on the signal to noise ratio (SNR) of the model parameters giving rise to benign versus non-benign, or harmful, overfitting: in particular, if the SNR is high then benign overfitting occurs, conversely if the SNR is low then harmful overfitting occurs. We attribute both benign and non-benign overfitting to an approximate margin maximization property and show that leaky ReLU networks trained on hinge loss with Gradient Descent (GD) satisfy this property. In contrast to prior work we do not require near orthogonality conditions on the training data: notably, for input dimension $d$ and training sample size $n$, while prior work shows asymptotically optimal error when $d=\Omega(n^{2}\log n)$, here we require only $d=\Omega\left(n\log\frac{1}{{\epsilon}}\right)$ to obtain error within ${\epsilon}$ of optimal.
@@ 1 Introduction
Intuition from learning theory suggests that fitting noise during training reduces a model’s performance on test data. However, it has been observed in some settings that machine learning models can interpolate noisy training data with only nominal cost to their generalization performance (Zhang et al., 2017; Belkin et al., 2018, 2019), a phenomenon referred to as benign overfitting. Establishing theory which can explain this phenomenon has attracted much interest in recent years and there is now a rich body of work on this topic particularly in the context of linear models. However, the study of benign overfitting in the context of non-linear models, in particular shallow ReLU or leaky ReLU networks, has additional technical challenges and subsequently is less well advanced.
@@
Much of the effort in regard to theoretically characterizing benign overfitting focuses on showing, under an appropriate scaling of the dimension of the input domain $d$, size of the training sample $n$, number of corruptions $k$ and number of model parameters $p$ that a model can interpolate noisy training data while achieving an arbitrarily small generalization error. Such characterizations of benign overfitting position it as a high dimensional phenomenon [^1] We provide a formal definition of benign overfitting as a high dimensional phenomenon in Appendix E.: indeed, the decrease in generalization error is achieved by escaping to higher dimensions at some rate relative to the other aforementioned hyperparameters. However, for these mathematical results to be relevant for explaining benign overfitting as observed in practice, clearly the particular scaling of $d$ with respect to $n,k$ and $p$ needs to reflect the ratios seen in practice. Although a number of works, which we discuss in Section 1.2, establish benign overfitting results for shallow neural networks, a key and significant limitation they share is the requirement that the input features of the training data are at least approximately orthogonal to one another. To study benign overfitting these prior works typically assume the input features consist of a small, low-rank signal component plus an isotropic noise term. Therefore, for the near orthogonality property to hold with high probability it is required that the input dimension $d$ scales as $d=\Omega(n^{2}\log n)$ or higher. This assumption highly restricts the applicability of these results for explaining benign overfitting in practice.
@@
In this work we assume only $d\approx\Omega(n)$ and establish both harmful and benign overfitting results for shallow leaky ReLU networks trained via Gradient Descent (GD) on the hinge loss. In particular, we consider $n$ data point pairs $({\bm{x}}_{i},y_{i})\in\mathbb{R}^{d}\times\{\pm 1\}$, where, for some vector ${\bm{v}}\in\mathbb{S}^{d-1}$ and scalar $\gamma\in[0,1]$, the input features are drawn from a pair of Gaussian clusters ${\bm{x}}_{i}\sim\mathcal{N}(\pm\sqrt{\gamma}{\bm{v}},\sqrt{1-\gamma}\tfrac{1}{d}(\textbf{I}_{d}-{\bm{v}}{\bm{v}}^{T}))$ and $y_{i}=\operatorname{sign}(\langle\mathbb{E}[{\bm{x}}_{i}],{\bm{v}}\rangle)$. The training data is noisy in that $k$ of the $n$ points in the training sample have their output label flipped. We assume equal numbers of positive and negative points among clean and corrupt ones. We provide a full description of our setup and assumptions in Section 2. Our proof techniques are novel and identify a new condition allowing for the analysis of benign and harmful overfitting which we term approximate margin maximization, wherein the norm of the network parameters is upper bounded by a constant of the norm of the max-margin linear classifier.
@@ 1.1 Summary of contributions
  • In Theorem 3.2, we establish conditions under which benign overfitting occurs for leaky ReLU networks. If the input dimension $d$, number of training points $n$, and signal strength $\gamma$ satisfy $d=\Omega(n)$ and $\gamma=\Omega(\frac{1}{n})$, then the network will exhibit benign overfitting. We emphasize that existing works on benign overfitting require $d=\Omega(n^{2}\log n)$ to ensure nearly orthogonal data.
@@ 1.2 Related work
There is now a significant body of literature theoretically characterizing benign overfitting in the context of linear models, including linear regression (Bartlett et al., 2020; Muthukumar et al., 2020; Wu & Xu, 2020; Zou et al., 2021; Hastie et al., 2022; Koehler et al., 2021; Wang et al., 2021a; Chatterji & Long, 2022; Shamir, 2022), logistic regression (Chatterji & Long, 2021; Muthukumar et al., 2021; Wang et al., 2021b), max-margin classification with linear and random feature models (Montanari et al., 2023b, a; Mei & Montanari, 2022; Cao et al., 2021) and kernel regression (Liang & Rakhlin, 2020; Liang et al., 2020; Adlam & Pennington, 2020). However, the study of benign overfitting in non-linear models is more nascent.
@@
Homogeneous networks trained with gradient descent and an exponentially tailed loss are known to converge in direction to a Karush-Kuhn-Tucker (KKT) point of the associated max-margin problem (Lyu & Li, 2020; Ji & Telgarsky, 2020) [^2] One also needs to assume initialization from a position with a low initial loss.. This property has been widely used in prior works to prove benign overfitting results for shallow neural networks. Frei et al. (2022) consider a shallow, smooth leaky ReLU network trained with an exponentially tailed loss and assume the data is drawn from a mixture of well-separated sub-Gaussian distributions. A key result of this work is, given sufficient iterations of GD, that the network will interpolate noisy training data while also achieving minimax optimal generalization error up to constants in the exponents. Xu & Gu (2023) extend this result to more general activation functions, including ReLU, as well as relax the assumptions on the noise distribution to being centered with bounded logarithmic Sobolev constant, and finally also improve the convergence rate. George et al. (2023) also study ReLU as opposed to leaky ReLU networks but do so in the context of the hinge loss, for which, and unlike exponentially tailed losses, a characterization of the implicit bias is not known. This work also establishes transitions on the margin of the clean data driving harmful, benign and no-overfitting training outcomes. Frei et al. (2023) use the aforementioned implicit bias of GD for linear classifiers and shallow leaky ReLU networks towards solutions which satisfy the KKT conditions of the margin maximization problem to establish settings where the satisfaction of said KKT conditions implies benign overfitting. Kornowski et al. (2023) also use the implicit bias results for exponentially tailed losses to derive similar benign overfitting results for shallow ReLU networks. Cao et al. (2022); Kou et al. (2023) study benign overfitting in two-layer convolutional as opposed to feedforward neural networks: indeed, whereas in most prior works data is modeled as the sum of a signal and noise component, in these two works the signal and noise components are assumed to lie in disjoint patches. The weight vector of each neuron is applied to both patches separately and a non-linearity, such as ReLU, is applied to the resulting pre-activation. In this setting, the authors prove interpolation of the noisy training data and derive conditions on the clean margin under which the network benignly versus harmfully overfits. A follow up work (Chen et al., 2023) considers the impact of Sharpness Aware Minimization (SAM) in the same setting. Finally, and assuming $d=\Omega(n^{5})$, Xu et al. (2024) establish benign overfitting results for a data distribution which, instead of being linearly separable, is separated according to an XOR function.
@@ 2.1 Data model
    ${\bm{y}}\in\{\pm 1\}^{n}$ is a random vector with entries $y_{i}$ which are mutually independent of one another as well as the noise vectors $({\bm{n}}_{i})_{i\in[n]}$ and are uniformly distributed over {±$\{\pm 1\}$. This vector holds the true labels of the training set.
@@
For finite $n$, this condition is equivalent to the existence of a ${\bm{w}}$ with $\hat{y}_{i}\langle{\bm{w}},{\bm{x}}_{i}\rangle>0$ for all $i\in[n]$. We denote the set of linearly separable datasets as $\mathcal{X}_{lin}\subset\mathbb{R}^{n\times d}\times\{\pm 1\}^{n}$. For a linearly separable dataset $({\bm{X}},\hat{{\bm{y}}})$, the max-margin linear classifier is the unique solution to the optimization problem
@@
Observe one may equivalently take a strictly convex objective $\|{\bm{w}}\|^{2}$ and the constraint set is a closed convex polyhedron which is non-empty iff the data is linearly separable. The max-margin linear classifier ${\bm{w}}^{\ast}$ has a corresponding geometric margin $2/\|{\bm{w}}^{\ast}\|$. When $d\geq n$ and $\gamma>0$, input feature matrices ${\bm{X}}$ from our data model almost surely have linearly independent rows ${\bm{x}}_{i}$ and thus $({\bm{X}},\hat{{\bm{y}}})$ is almost surely linearly separable for any observed labels $\hat{{\bm{y}}}\in\{\pm 1\}^{n}$.
@@ 2.2 Architecture and learning algorithm
where ${\bm{W}}\in\mathbb{R}^{2m\times d}$ are the parameters of the network, $\sigma:\mathbb{R}\to\mathbb{R}$ is the leaky ReLU function, defined as $\sigma(x)=\max(x,\alpha x)$, where $\alpha\in(0,1]$ is referred to as the leaky parameter. We remark that we only train the weights of the first layer and keep the output weights of each neuron fixed. Although $\sigma$ is not differentiable at 00, in the context of gradient descent we adopt a subgradient and let $\dot{\sigma}(z)=1$ for $z\geq 0$ and let $\dot{\sigma}(z)=\alpha$ otherwise. The hinge loss $\ell:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ is defined as
@@
The step size $\eta$ satisfies $\eta\leq 1/(mn\max_{i\in[n]}\|{\bm{x}}_{i}\|^{2})$ and for all $j\in[2m]$ the network at initialization satisfies $\|{\bm{w}}_{j}^{(0)}\|\leq\sqrt{\alpha}/(m\min_{i\in[n]}\|{\bm{x}}_{i}\|)$.
@@ 3 Main results
In order to prove benign overfitting it is necessary to show that the learning algorithm outputs a model which correctly classifies all points in the training sample. The following theorem establishes even stronger conditions for $\mathcal{A}_{GD}$.
@@
Furthermore $\mathcal{A}_{GD}$ is approximately margin maximizing on $f$ with
@@
  • $\mathcal{A}:\mathbb{R}^{n\times d}\times\{\pm 1\}^{n}\rightarrow\mathbb{R}^{2m\times d}$ is a learning algorithm which returns the weights ${\bm{W}}\in\mathbb{R}^{2m\times d}$ of the first layer of $f$.
@@
Under the setting given in Definition 3.1, let $\delta,\epsilon\in(0,1)$ and suppose $\mathcal{A}$ is approximately margin-maximizing. If $n=\Omega\left(\log\frac{1}{\delta}\right)$, $d=\Omega\left((1+|\mathcal{A}|^{2}m)n\log\frac{1}{\epsilon}\right)$, $k=O(\frac{n}{1+m|\mathcal{A}|^{2}})$, and $\gamma=\Omega\left(\frac{1+m|\mathcal{A}|^{2}}{n}\right)$ then with probability at least $1-\delta$ over $({\bm{X}},\hat{{\bm{y}}})$
$\mathbb{P}(yf({\bm{W}},{\bm{x}})\leq 0\mid{\bm{X}},\hat{{\bm{y}}})\leq\epsilon.$
A proof of Theorem 3.2 is provided in Appendix D.2. Combining Theorems 3.1 and 3.2 we arrive at the following benign overfitting result for shallow leaky ReLU networks trained with GD on hinge loss.
@@
Under the setting given in Definition 3.1, let $\delta,\epsilon\in(0,1)$ and suppose $\mathcal{A}=\mathcal{A}_{GD}$ where $\eta,\lambda\in\mathbb{R}_{>0}$ satisfy Assumption 1. If $n=\Omega\left(\log\frac{1}{\delta}\right)$, $d=\Omega\left(n\log\frac{1}{\epsilon}\right)$, $k=O(n)$, and $\gamma=\Omega\left(\frac{1}{n}\right)$ then the following hold.
@@
    Almost surely $\mathcal{A}_{GD}$ terminates after a finite number of updates and with ${\bm{W}}=\mathcal{A}_{GD}({\bm{X}},\hat{{\bm{y}}})$ then $L({\bm{W}},{\bm{X}},\hat{{\bm{y}}})=0$.
@@
    $\mathbb{P}(yf({\bm{W}},{\bm{x}})\leq 0\mid{\bm{X}},\hat{{\bm{y}}})\leq\epsilon.$
@@
Under the setting given in Definition 3.1, let $\delta\in(0,1)$ and suppose $\mathcal{A}=\mathcal{A}_{GD}$, where $\eta,\lambda\in\mathbb{R}_{>0}$ satisfy Assumption 1. If $n=\Omega(1),d=\Omega\left(n+\log\frac{1}{\delta}\right)$ and $\gamma=O\left(\frac{\alpha^{3}}{d}\right)$ then the following hold.

### v2

@@ Abstract
The problem of benign overfitting asks whether it is possible for a model to perfectly fit noisy training data and still generalize well. We study benign overfitting in two-layer leaky ReLU networks trained with the hinge loss on a binary classification task. We consider input data that can be decomposed into the sum of a common signal and a random noise component, that lie on subspaces orthogonal to one another. We characterize conditions on the signal to noise ratio (SNR) of the model parameters giving rise to benign versus non-benign (or harmful) overfitting: in particular, if the SNR is high then benign overfitting occurs, conversely if the SNR is low then harmful overfitting occurs. We attribute both benign and non-benign overfitting to an approximate margin maximization property and show that leaky ReLU networks trained on hinge loss with gradient descent (GD) satisfy this property. In contrast to prior work we do not require the training data to be nearly orthogonal. Notably, for input dimension $d$ and training sample size $n$, while results in prior work require $d=\Omega(n^{2}\log n)$, here we require only $d=\Omega\left(n\right)$.
@@ 1 Introduction
Intuition from learning theory suggests that fitting noise during training reduces a model’s performance on test data. However, it has been observed in some settings that machine learning models can interpolate noisy training data with only nominal cost to their generalization performance (Zhang et al., 2017; Belkin et al., 2018, 2019),
a phenomenon referred to as benign overfitting. Establishing theory that can explain this phenomenon has attracted much interest in recent years and there is now a rich body of work on this topic particularly in the context of linear models. However, the study of benign overfitting in the context of non-linear models, in particular shallow ReLU or leaky ReLU networks, has additional technical challenges and subsequently is less well advanced.
@@
Much of the effort in regard to theoretically characterizing benign overfitting focuses on showing, under an appropriate scaling of the dimension of the input domain $d$, size of the training sample $n$, number of corruptions $k$ and number of model parameters $p$ that a model can interpolate noisy training data while achieving an arbitrarily small generalization error. Such characterizations of benign overfitting position it as a high dimensional phenomenon [^1] We provide a formal definition of benign overfitting as a high dimensional phenomenon in Appendix E.: indeed, the decrease in generalization error is achieved by escaping to higher dimensions at some rate relative to the other aforementioned hyperparameters. However, for these mathematical results to be relevant for explaining benign overfitting as observed in practice, clearly the particular scaling of $d$ with respect to $n,k$ and $p$ needs to reflect the ratios seen in practice. Although a number of works, which we discuss in Section 1.2, establish benign overfitting results for shallow neural networks, a key and significant limitation they share is the requirement that the input features of the training data are at least approximately orthogonal to one another. To study benign overfitting, these prior works typically assume the input features consist of a small, low-rank signal component plus an isotropic noise term. Therefore, for the near orthogonality property to hold with high probability it is required that the input dimension $d$ scales as $d=\Omega(n^{2}\log n)$ or higher. This assumption highly restricts the applicability of these results for explaining benign overfitting in practice.
@@
In this work we assume only $d=\Omega(n)$ and establish both harmful and benign overfitting results for shallow leaky ReLU networks trained via gradient descent (GD) on the hinge loss. In particular, we consider $n$ data point pairs $({\bm{x}}_{i},y_{i})\in\mathbb{R}^{d}\times\{\pm 1\}$, where, for some vector ${\bm{v}}\in\mathbb{S}^{d-1}$ and scalar $\gamma\in[0,1]$, the input features are drawn from a pair of Gaussian clusters ${\bm{x}}_{i}\sim\mathcal{N}(\pm\sqrt{\gamma}{\bm{v}},\sqrt{1-\gamma}\tfrac{1}{d}(\textbf{I}_{d}-{\bm{v}}{\bm{v}}^{T}))$ and $y_{i}=\operatorname{sign}(\langle\mathbb{E}[{\bm{x}}_{i}],{\bm{v}}\rangle)$. The training data is noisy in that $k$ of the $n$ points in the training sample have their output label flipped. We
assume equal numbers of positive and negative points among clean and corrupt ones. We provide a full description of our setup and assumptions in Section 2. Our proof techniques are novel and identify a new condition allowing for the analysis of benign and harmful overfitting which we term approximate margin maximization, wherein the norm of the network parameters is upper bounded by a constant of the norm of the max-margin linear classifier.
@@ 1.1 Summary of contributions
  • In Theorem 3.2, we establish conditions under which benign overfitting occurs for leaky ReLU networks. If the input dimension $d$, number of training points $n$, number of corrupt points $k$, and signal strength $\gamma$ satisfy $d=\Omega(n)$ and $\gamma=\Omega(\frac{1}{k})$, then the network will exhibit benign overfitting. We emphasize that existing works on benign overfitting require $d=\Omega(n^{2}\log n)$ to ensure nearly orthogonal data.
  • In Theorem 3.3, we find a complementary lower bound for the generalization error to show that, for gradient descent classifiers, the bound in Theorem 3.2 is tight up to a constant in the exponent that can depend on $\alpha$.
@@ 1.2 Related work
There is now a significant body of literature theoretically characterizing benign overfitting in the context of linear models, including linear regression (Bartlett et al., 2020; Muthukumar et al., 2020; Wu & Xu, 2020; Zou et al., 2021; Hastie et al., 2022; Koehler et al., 2021; Wang et al., 2021a; Chatterji & Long, 2022; Shamir, 2022),
logistic regression (Chatterji & Long, 2021; Muthukumar et al., 2021; Wang et al., 2021b), max-margin classification with linear and random feature models (Montanari et al., 2023b, a; Mei & Montanari, 2022; Cao et al., 2021) and kernel regression (Liang & Rakhlin, 2020; Liang et al., 2020; Adlam & Pennington, 2020).
However, the study of benign overfitting in non-linear models is more nascent.
@@
Homogeneous networks trained with gradient descent and an exponentially tailed loss are known to converge in direction to a Karush-Kuhn-Tucker (KKT) point of the associated max-margin problem (Lyu & Li, 2020; Ji & Telgarsky, 2020) [^2] One also needs to assume initialization from a position with a low initial loss.. This property has been widely used in prior works to prove benign overfitting results for shallow neural networks. Frei et al. (2022) consider a shallow, smooth leaky ReLU network trained with an exponentially tailed loss and assume the data is drawn from a mixture of well-separated sub-Gaussian distributions. A key result of this work is, given sufficient iterations of GD, that the network will interpolate noisy training data while also achieving minimax optimal generalization error up to constants in the exponents. Xu & Gu (2023) extend this result to more general activation functions, including ReLU, as well as relax the assumptions on the noise distribution to being centered with bounded logarithmic Sobolev constant, and finally also improve the convergence rate. George et al. (2023) also study ReLU as opposed to leaky ReLU networks but do so in the context of the hinge loss, for which, and unlike exponentially tailed losses, a characterization of the implicit bias is not known. This work also establishes transitions on the margin of the clean data driving harmful, benign and no-overfitting training outcomes. Frei et al. (2023) use the aforementioned implicit bias of GD for linear classifiers and shallow leaky ReLU networks towards solutions that satisfy the KKT conditions of the margin maximization problem to establish settings where the satisfaction of said KKT conditions implies benign overfitting. Kornowski et al. (2023) also use the implicit bias results for exponentially tailed losses to derive similar benign overfitting results for shallow ReLU networks. Cao et al. (2022); Kou et al. (2023) study benign overfitting in two-layer convolutional as opposed to feedforward neural networks: indeed, whereas in most prior works data is modeled as the sum of a signal and noise component, in these two works the signal and noise components are assumed to lie in disjoint patches. The weight vector of each neuron is applied to both patches separately and a non-linearity, such as ReLU, is applied to the resulting pre-activation. In this setting, the authors prove interpolation of the noisy training data and derive conditions on the clean margin under which the network benignly versus harmfully overfits. A follow up work (Chen et al., 2023) considers the impact of Sharpness Aware Minimization (SAM) in the same setting. Finally, and assuming $d=\Omega(n^{5})$, Xu et al. (2024) establish benign overfitting results for a data distribution which, instead of being linearly separable, is separated according to an XOR function.
@@ 2 Preliminaries
Let $[n]=\{1,2,\ldots,n\}$ denote the set of the first $n$ natural numbers.
We remark that when using big-$O$ notation we implicitly assume only positive constants. We use $c,C,C_{1},C_{2},\ldots$ to denote absolute constants with respect to the input dimension $d$, the training sample size $n$, and the width of the network $m$. Note constants may change in value from line to line. Furthermore, when using big-$O$ notation all variables aside from $d,n,k$ and $m$ are considered constants. However, for clarity we will frequently make the constants concerning the confidence $\delta$ and failure probability $\epsilon$ explicit. Moreover, for two functions $f,g:\operatorname{\mathbb{N}}\rightarrow\operatorname{\mathbb{N}}$, if we say $f=O(g)$ implies property $p$, what we mean is there exists an $N\in\operatorname{\mathbb{N}}$ and a constant $C$ such that if $f(n)\leq Cg(n)$ for all $n\geq N$ then property $p$ holds. Likewise, if we say $f=\Omega(g)$ implies property $p$, what we mean is there exists an $N\in\operatorname{\mathbb{N}}$ and a constant $C$
@@ 2.1 Data model
    ${\bm{y}}\in\{\pm 1\}^{n}$ is a random vector with entries $y_{i}$ that are mutually independent of one another as well as the noise vectors $({\bm{n}}_{i})_{i\in[n]}$ and are uniformly distributed over {±$\{\pm 1\}$. This vector holds the true labels of the training set.
@@
For finite $n$, this condition is equivalent to the existence of a ${\bm{w}}$ with $\hat{y}_{i}\langle{\bm{w}},{\bm{x}}_{i}\rangle>0$ for all $i\in[n]$.
We denote the set of linearly separable datasets as $\mathcal{X}_{lin}\subset\mathbb{R}^{n\times d}\times\{\pm 1\}^{n}$. For a linearly separable dataset $({\bm{X}},\hat{{\bm{y}}})$, the max-margin linear classifier is the unique solution to the optimization problem
@@
Observe one may equivalently take a strictly convex objective $\|{\bm{w}}\|^{2}$ and the constraint set is a closed convex polyhedron that is non-empty iff the data is linearly separable. The max-margin linear classifier ${\bm{w}}^{\ast}$ has a corresponding geometric margin $2/\|{\bm{w}}^{\ast}\|$. When $d\geq n$ and $\gamma>0$, input feature matrices ${\bm{X}}$ from our data model almost surely have linearly independent rows ${\bm{x}}_{i}$ and thus $({\bm{X}},\hat{{\bm{y}}})$ is almost surely linearly separable for any observed labels $\hat{{\bm{y}}}\in\{\pm 1\}^{n}$.
@@ 2.2 Architecture and learning algorithm
where ${\bm{W}}\in\mathbb{R}^{2m\times d}$ are the parameters of the network, $\sigma:\mathbb{R}\to\mathbb{R}$ is the leaky ReLU function, defined as $\sigma(x)=\max(x,\alpha x)$, where $\alpha\in(0,1]$ is referred to as the leaky parameter. We remark that we only train the weights of the first layer and keep the output weights of each neuron fixed. Although $\sigma$ is not differentiable at 00, in the context of gradient descent we adopt a subgradient and let $\dot{\sigma}(z)=1$ for $z\geq 0$ and let $\dot{\sigma}(z)=\alpha$ otherwise.
The hinge loss $\ell:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ is defined as
@@
The step size $\eta$ satisfies $\eta\leq 1/(mn\max_{i\in[n]}\|{\bm{x}}_{i}\|^{2})$
and for all $j\in[2m]$ the network at initialization satisfies $\|{\bm{w}}_{j}^{(0)}\|\leq\sqrt{\alpha}/(m\min_{i\in[n]}\|{\bm{x}}_{i}\|)$.
@@ 3 Main results
In order to prove benign overfitting it is necessary to show that the learning algorithm outputs a model that correctly classifies all points in the training sample. The following theorem establishes this for $\mathcal{A}_{GD}$ and bounds the margin maximizing factor $|\mathcal{A}_{GD}|$.
@@
Furthermore $\mathcal{A}_{GD}$ is approximately margin maximizing on $f$ (Definition 2.3) with
@@
  • $\mathcal{A}:\mathbb{R}^{n\times d}\times\{\pm 1\}^{n}\rightarrow\mathbb{R}^{2m\times d}$ is a learning algorithm that returns the weights ${\bm{W}}\in\mathbb{R}^{2m\times d}$ of the first layer of $f$.
@@
Under the setting given in Assumption 2, let $\delta\in(0,1)$ and suppose $\mathcal{A}$ is approximately margin-maximizing (Definition 2.3). If $n=\Omega\left(\log\frac{1}{\delta}\right)$, $d=\Omega\left(n\right)$, $k=O(\frac{n}{1+m|\mathcal{A}|^{2}})$, and $\gamma=\Omega\left(\frac{1}{k}\right)$ then there is a fixed positive constant $C$ such that with probability at least $1-\delta$ over $({\bm{X}},\hat{{\bm{y}}})$
$\mathbb{P}(yf({\bm{W}},{\bm{x}})\leq 0\mid{\bm{X}},\hat{{\bm{y}}})\leq\exp\left(-C\cdot\frac{d}{k(1+m|\mathcal{A}|^{2})}\right).$
A proof of Theorem 3.2 is provided in Appendix D.2. Combining Theorems 3.1, and 3.2 we arrive at the following benign overfitting result for shallow leaky ReLU networks trained with GD on hinge loss.
@@
Under the setting given in Assumption 2, let $\delta\in(0,1)$ and suppose $\mathcal{A}=\mathcal{A}_{GD}$ where $\eta,\lambda\in\mathbb{R}_{>0}$ satisfy Assumption 1. If $n=\Omega\left(\log\frac{1}{\delta}\right)$, $d=\Omega\left(n\right)$, $k=O(\alpha^{2}n)$, and $\gamma=\Omega\left(\frac{1}{k}\right)$ then the following hold.
@@
    The algorithm $\mathcal{A}_{GD}$ terminates almost surely after a finite number of updates. If ${\bm{W}}=\mathcal{A}_{GD}({\bm{X}},\hat{{\bm{y}}})$, then $L({\bm{W}},{\bm{X}},\hat{{\bm{y}}})=0$.
@@
    There is a fixed positive constant $C$ such that, with probability at least $1-\delta$ over the training data $({\bm{X}},\hat{{\bm{y}}})$,
    $\mathbb{P}(yf({\bm{W}},{\bm{x}})\leq 0\mid{\bm{X}},\hat{{\bm{y}}})\leq\exp\left(-C\cdot\frac{\alpha^{2}d}{k}\right).$
We remark that the upper bound is at most $\exp\left(-Cd/n\right)$ for a different constant $C$ as we assume $k=O(\alpha^{2}n)$.
If $k$ is large enough, this bound is tight up to constants and factors of $\alpha$ in the exponent. This is given by the following theorem, proven in Appendix D.2.
Theorem 3.3.
Under the setting given in Assumption 2, let $\delta\in(0,1)$ and suppose $\mathcal{A}=\mathcal{A}_{GD}$ where $\eta,\lambda\in\mathbb{R}_{>0}$ satisfy Assumption 1. If $n=\Omega\left(k\right)$, $d=\Omega\left(n\right)$, and $k=\Omega(\log\frac{1}{\delta}+\frac{1}{\alpha})$, then there is a fixed positive constant $C$ such that with probability at least $1-\delta$ over $({\bm{X}},\hat{{\bm{y}}})$
$\mathbb{P}(yf({\bm{W}},{\bm{x}})\leq 0\mid{\bm{X}},\hat{{\bm{y}}})\geq\exp\left(-C\cdot\frac{d}{\alpha k}\right).$
@@
Under the setting given in Assumption 2, let $\delta\in(0,1)$ and suppose $\mathcal{A}=\mathcal{A}_{GD}$, where $\eta,\lambda\in\mathbb{R}_{>0}$ satisfy Assumption 1. If $n=\Omega(1),d=\Omega\left(n+\log\frac{1}{\delta}\right)$ and $\gamma=O\left(\frac{\alpha^{3}}{d}\right)$ then the following hold.